Please help me solve this equation and show each step: (x+y)^2
Answer by norymae(15)

longer solution:..
use parrot beak:
(x+y)(x+y)
x(x)+x(y)+y(x)+y(y)
x^2+xy+xy+y^2
=x^2+2xy+y^2
shorter solution:..
(x+y)^2
x=1st term
y=2nd term
*square the first term =x^2
*double the product of the 2 terms =2(x*y)=2xy
*square the last term y^2
*then add
=x^2+2xy+y^2